Preparing a 10% w/v NaCl solution



Determine the amount in grams of NaCl needed to prepare 10% w/v NaCl



solution. Weigh this amount of sodium chloride in a 50 mL beaker. Add about 30



mL distilled water and stir. Once the sodium chloride has dissolved completely, pour



it into a 100 mL volumetric flask. Add distilled water up to the mark to bring the



volume to 100 ml. Mix thoroughly. Caution: Do not simply measure 100ml of water



and add 10g of sodium chloride. This will introduce error because adding the solid



will change the final volume of the solution and throw off the final percentage




Preparing a 10% w/v NaCl solution



Grams of NaCl = _______________



Calculation:



1
Expert's answer
2021-12-06T03:35:02-0500

10%w/v = 0.10

Mass in grams = 0.10 × 50 = 5grams